OpenStudy (anonymous):

3x^2+2x-1=0 is it imaginary irrational equal rational

jimthompson5910 (jim_thompson5910):

use the discriminant formula D = b^2 - 4ac to figure out how many and what type of solutions there are

OpenStudy (anonymous):

i do not kn ow that formula

jimthompson5910 (jim_thompson5910):

3x^2+2x-1 is in the form ax^2+bx+c

jimthompson5910 (jim_thompson5910):

in this case a = 3 b = 2 c = -1

jimthompson5910 (jim_thompson5910):

plug all that into D = b^2 - 4ac and tell me what you get

OpenStudy (anonymous):

17

jimthompson5910 (jim_thompson5910):

D = b^2 - 4ac D = (2)^2 - 4(3)(-1) D = 4 + 12 D =16

OpenStudy (anonymous):

D=17

jimthompson5910 (jim_thompson5910):

since this is a) positive and b) a perfect square this means that you have 2 different rational solutions

jimthompson5910 (jim_thompson5910):

how are you getting 17

OpenStudy (anonymous):

disc=2^2-4*3*-1=16=4^2, which is positive and a perfect square Hence it is real and rational .

OpenStudy (anonymous):

roots are equal only when D=0

OpenStudy (anonymous):

i got 17 because i put the problem in right. you used two i used three for a.

jimthompson5910 (jim_thompson5910):

no I used a = 3, not a = 2

jimthompson5910 (jim_thompson5910):

I used b = 2, you must be mixing those up

OpenStudy (anonymous):

my bad i wrote the formula wrong hahaha ok i got 16 now.

OpenStudy (anonymous):

D>0 and a perfect square hence roots are real,rational and different.

OpenStudy (anonymous):

 can you tell me how to work this i have no idea where to start